NON-PLANAR STRUCTURES OF THE HIGH-ENERGY ROTATIONAL CONFORMERS OF 2-METHYLBUTA-1,3-DIENE (ISOPRENE) AND 2,3-DIMETHYLBUTA-1,3-DIENE

\underline{Yu. N. PANCHENKO}, Dept. of Chemistry, M. V. Lomonosov Moscow State University, Moscow 119992, Russian Federation; Ch. W. BOCK, Dept. of Chemistry and Biochemistry, Philadelphia University, Philadelphia, PA 19144; J. D. LARKIN, Dept. of Chemistry, Bloomsburg University of Pennsylvania, Bloomsburg, PA 17815; A. V. ABRAMENKOV, Dept. of Chemistry, M. V. Lomonosov Moscow State University, Moscow 119992, Russian Federation; and F. KÜHNEMANN, Institute of Applied Physics, Bonn University, 53115 Bonn, Germany.

Optimization of the geometrical parameters and determination of the force fields for rotamers of the title molecules were performed at the MP2(FC)/aug-cc-pVDZ//MP2(FC)/aug-cc-pVDZ computational level. The vibrational analyses of these conformers were carried out using scaled quantum-mechanical force field methodology. Recent experimental wavenumbers for these conformers and their deuteroisomers were incorporated into these analyses. The theoretical non-planar structures of the high-energy conformers of 2-methylbuta-1,3-diene (isoprene) and 2,3-dimethylbuta-1,3-diene were corroborated by good agreement between the experimental and theoretical wavenumbers of the molecules under investigation. The dihedral angles of the non-planar high-energy conformers for rotation around the =C–C= bond are as follows: 41.6^{∘} for 2-methylbuta-1,3-diene (isoprene)^{a} and 47.0^{∘} for 2,3-dimethylbuta-1,3-diene^{b}. Previous studies performed at the HF/6-31G level gave 41.0^{∘} and 48.5^{∘} for the first and second compounds, respectively.

aYu. N. Panchenko, Ch. W. Bock, J. D. Larkin, A. V. Abramenkov, F. Kühnemann, Struct. Chem. \underline{19}, 421 (2008).

bYu. N. Panchenko, Ch. W. Bock, J. D. Larkin, A. V. Abramenkov, Struct. Chem. \underline{19}, 793 (2008).